# The Holographic QCD Running Coupling Constant from the Ricci Flow

### v1

@@ 4 Masses and decay constants
The proposed dilaton Eq.(23) can be used to predict the vector mesons Regge trajectory after solving their corresponding equation of motion [6]:
@@
[…]
0 775 346
1 1154 393
2 1427 435
3 1652 469
4 1850 496
5 2028 519
6 2192 540
7 2344 558
8 2487 575
Table 2: Masses and decay constants of vector mesons predicted by the DDSW model Eq.(23).
Regge trajectories for our DDSW model are linear as in the Soft Wall model, that is, $M_{n}^{2}\sim n$. We depict our predictions using the DDSW model for the $\rho$-meson family (red points in Fig.3) together with the physical values from the PDG [28] (purple boxes) where errors are obtained via de half-width rule (i.e, $M_{n}^{2}\pm M_{n}\Gamma_{n}$, with $\Gamma_{n}$ the total decay width of the particle [29]), and the prediction of Soft Wall model [6] (blue stars). Comparison to experimental data demands merging all $\rho$ states with total angular momentum $J=1$, including 1S and 2S states, as we understand the holographic model cannot distinguish among them. In this respect, the Regge slope for the radial trajectory for the 7 known members of the $\rho$ family would read $0.87(7)$GeV-2 as an average of the 1S and 2S trajectories from Ref.[30]. The DDSW model yields a slope $0.6962(18)$ GeV-2.
@@
Figure 3: Regge trajectories $(M_{n}^{2},n)$ for the DDSW model (red points), the SW model (blue stars), and comparison with experimental data (purple boxes) [28].
@@ 5 Conclusions
In this work, we have explored a new holographic model of QCD, named Double Dilaton Soft Wall model (DDSW), which from the Ricci flow allow us to obtain a parameterization for $\alpha_{s}$ able to match an infrared fixed point at low energies with pQCD at high energies. We used an AdS/CFT dictionary formula defined in Eq.(3), to build the DDSW raising from a dilaton background defined in Eq.(23) that breaks conformal symmetry geometrically thanks to including two dilaton fields of opposite signs (cf. [18, 19, 20]), each of them understood as vector and axial-vector mesons seeing different geometries [17]. The matching with pQCD, which is found way above 2 GeV, is successful thanks to including in the holographic side large-$N_{c}$ corrections. In the process, we use experimental data from Ref. [23] to fit the free leftover parameters of the $\alpha_{s}$ parameterization.
As a by-product, we have studied the spectrum of vector mesons in the DDSW model by solving the equation of motion of the vector gauge field [6] with the dilaton mentioned above. This allowed us to determine masses as eigenvalues of the differential operator defined by the equation of motion, obtaining linear Regge trajectories. The predicted masses are heavier than in the usual Soft Wall model, thus closer to experimental results. Moreover, the squared root of the decay constant of the $\rho$-meson is improved compared to usual SW models as well, giving a departure of less than $1\%$ when compared to the experimental result.
@@ Acknowledgements
The work of PM has been supported by the Ministerio de Ciencia e Innovación under grant PID2020-112965GB-I00 and by the Secretaria d’Universitats i Recerca del Departament d’Empresa i Coneixement de la Generalitat de Catalunya under grant 2021 SGR 00649. IFAE is partially funded by the CERCA program of the Generalitat de Catalunya.

### v2

@@ 4 Masses and decay constants
The proposed dilaton Eq.(23) can be used to predict different meson Regge trajectory after solving their corresponding equation of motion [6]. For vector mesons the five dimensional mass is equal to zero (in general it is given by $M_{5}^{2}=(\Delta-S)(\Delta+S-4)$, where $\Delta$ is the dimension of the operator dual to the field in QCD and $S$ is the spin). The corresponding equation of motion is:
@@
$n$ Scalar $M_{n}\text{ (MeV)}$ Vector $M_{n}\text{ (MeV)}$ Tensor $M_{n}\text{ (MeV)}$
0 944 775 1144
1 1281 1154 $14$
2 1534 1427 1642
3 1749 1652 1840
4 1940 1850 2017
5 2114 2028 2180
6 2276 2192 2330
7 2428 2344 2471
8 2571 2487 2605
Table 2: Masses of scalar, vector and tensor mesons predicted by the DDSW model Eq.(23).
Regge trajectories for our DDSW model are linear as in the Soft Wall model, that is, $M_{n}^{2}\sim n$. We depict our predictions using the DDSW model for the $\rho$-meson family (red points in Fig.3) together with the physical values from the PDG [32] (purple boxes) where errors are obtained via de half-width rule (i.e, $M_{n}^{2}\pm M_{n}\Gamma_{n}$, with $\Gamma_{n}$ the total decay width of the particle [33]), and the prediction of Soft Wall model [6] (blue stars). Comparison to experimental data demands merging all $\rho$ states with total angular momentum $J=1$, including 1S and 2S states, as we understand the holographic model cannot distinguish among them. In this respect, the Regge slope for the radial trajectory for the 7 known members of the $\rho$ family would read 0.87(7)0.87(7)GeV-2 as an average of the 1S and 2S trajectories from Ref.[34]. The DDSW model yields a slope $0.6962(18)$ GeV-2.
@@
Figure 3: Vector meson Regge trajectories $(M_{n}^{2},n)$ for the DDSW model (red points), the SW model (blue stars), and comparison with experimental data (purple boxes) [32], [43].
@@
to obtain $\Gamma(\rho^{0}\to e^{+}e^{-})=6.87\text{ keV}$. This gives a departure of $2.41\%$ in comparison to the experimental value $7.04(6)\text{ keV}$ [32]. For the usual Soft Wall model the value is $2.19\text{ keV}$ with a departure of $68.9\%$.
Using the same procedure we can find the spectrum of scalar and tensor mesons. Scalar mesons are particularly interesting since they have a non zero five dimensional mass $M_{5}^{2}=-3$. Following the usual scalar field action [44], we must solve:
@@
$=-M_{n}^{2}\frac{e^{-\phi^{\Delta}(z)}}{z^{3}}\psi_{n}(z),$
with the normalization:
$\int_{0}^{+\infty}dz\frac{e^{-\phi^{\Delta}(z)}}{z^{3}}|\psi_{n}(z)|^{2}=1.$ (40)
Results are depicted in the second column of Table 2, in which an improvement respect the original Soft Wall model is achieved, and more accurate results are obtained when comparing to experimental data [43]. In fact, resonances with $n=2,3,5$ show a departure to the experimental counterpart below $1\%$. In Figure 4, the empty circle and the empty star represent the prediction of DDSW and Soft Wall models, respectively, for $f_{0}(2330)$ meson. We have not included this resonance as a experimental data point since there is no consensus for its mean value [43]. In this way we predict the mass of the $f_{0}(2330)$ as $2428\text{ MeV}$ in the DDSW model. We observe this value is similar to that of [45], which gives a mass of $2419\pm 64\text{ MeV}$, heavier than other analysis present in the PDG [43].
Figure 4: Scalar meson Regge trajectories $(M_{n}^{2},n)$ for the DDSW model (red points), the SW model (blue stars), and comparison with experimental data (purple boxes) [43].
$n$ Scalar $F_{n}^{1/2}\text{ (MeV)}$ Vector $F_{n}^{1/2}\text{ (MeV)}$ Tensor $F_{n}^{1/2}\text{ (MeV)}$
0 289 346 0.852
1 317 393 106
2 346 435 122
3 371 469 137
4 390 496 150
5 408 519 163
6 423 540 175
7 438 558 186
8 457 575 202
Table 4: Decay constants of scalar, vector and tensor mesons predicted by the DDSW model.
Finally we adress the spectrum of tensor mesons. Here $M_{5}^{2}=0$, so the corresponding equation of motion is:
$\partial_{z}\left(\frac{e^{-\phi^{\Delta}(z)}}{z^{3}}\partial_{z}\psi_{n}(z)\right)=-M_{n}^{2}\frac{e^{-\phi^{\Delta}(z)}}{z^{3}}\psi_{n}(z),$ (41)
with the same normalization as in the scalar case Eq.(40). Some studies of tensor mesons distinguish four different Regge trajectories [46], [47]. We have introduced tensor mesons with the same action as in SW model [6], [48] introducing our dilaton field. The output of our model is a single trajectory, so we have compared our predicted masses with the experimental data available. More precisely, our model is a large-$N_{c}$ model, so we have stable mesons, with zero decay widths. As consequence, we have grouped the resonances found in experiment if they are overlaped by its decay width. For example, $f_{2}(1950)$ has a large decay width ($464\pm 24\text{ MeV}$, [43]) that overlaps with other resonances like $f_{2}(2010)$. In this case we have plotted only one resonance, the $f_{2}(2010)$, instead of two, by comparing with our predictions from DDSW model. In this way, the Regge trajectory in Table 5 is a proposal of organising the $f_{2}$ resonances.
As a summary of results, we conclude that thanks to including not only the $z^{2}$ behavior at large $z$ -which renders the Regge trajectories linear- but also the $z\to 0$ constant behavior of DDSW dilaton, our mass predictions are closer to their experimental counterparts compared to the original SW model.
Figure 5: Proposal for tensor meson Regge trajectory $(M_{n}^{2},n)$ for the DDSW model (red points), the SW model (blue stars), and comparison with experimental data (purple boxes)[43].
@@ 5 Conclusions
In this work, we have explored a new holographic model of QCD, named Double Dilaton Soft Wall model (DDSW), which from the Ricci flow allow us to obtain a parameterization for $\alpha_{s}$ able to match an infrared fixed point at low energies with pQCD at high energies. We used an AdS/CFT dictionary formula defined in Eq.(3), to build the DDSW raising from a dilaton background defined in Eq.(23) that breaks chiral symmetry geometrically thanks to including two dilaton fields of opposite signs (cf. [22, 23, 24]), each of them understood as vector and axial-vector mesons seeing different geometries [21]. The matching with pQCD, which is found way above 2 GeV, is successful thanks to including in the holographic side large-$N_{c}$ corrections. In the process, we use experimental data from Ref. [27] to fit the free leftover parameters of the $\alpha_{s}$ parameterization. Because our model Eq.(31) have various adjustable parameters which limits its predictive power, we expect to improve it in the future.
As a by-product, we have studied the spectrum of vector mesons in the DDSW model by solving the equation of motion [6] of different meson families with the dilaton mentioned above. This allowed us to determine masses as eigenvalues of the differential operator defined by the equation of motion, obtaining linear Regge trajectories for scalar, vector and tensor mesons. In particular, the predicted scalar and vector meson masses are heavier than in the usual Soft Wall model, thus closer to experimental results. Based on these results, we have done a proposal for the tensor meson Regge trajectory. Moreover, the squared root of the decay constant of the $\rho$-meson is improved compared to usual SW models, giving a departure of less than $1\%$ compared to the experimental result.
@@ Acknowledgements
We are grateful to V. Vento for discussions and encouregement.
The work of PM has been supported by the Ministerio de Ciencia e Innovación under grant PID2020-112965GB-I00 and by the Secretaria d’Universitats i Recerca del Departament d’Empresa i Coneixement de la Generalitat de Catalunya under grant 2021 SGR 00649. The CERCA program of the Generalitat de Catalunya partially funds IFAE.